OpenStudy (konradzuse):

find the integral (cos(x))^9 sin(x)dx

OpenStudy (konradzuse):

ok so...

OpenStudy (turingtest):

and what exactly is the problem here? I've seen you do much harder than this

OpenStudy (konradzuse):

\[\sin(x)(\cos(x)^2)^4 \cos(x)\]

OpenStudy (turingtest):

why?

OpenStudy (turingtest):

why would you want to do that?

OpenStudy (konradzuse):

Just going by the section stuff :P.

OpenStudy (konradzuse):

Maybe it's easier than I'm thinking?

OpenStudy (turingtest):

also\[\cos(x)^2\]is very confusing notation, does it mean\[\cos^2x\]or\[\cos(x^2)\]

OpenStudy (turingtest):

and yes, it's way easier than you are thinking

OpenStudy (unklerhaukus):

standard table of integrals for \[\int\sin ^nx\cos^mx\text dx\]

OpenStudy (konradzuse):

I guess (cos(x))^9 is the same as cos^9(x)?

OpenStudy (turingtest):

right

OpenStudy (konradzuse):

what rhauskus said :P

OpenStudy (turingtest):

but that is for the general case your problem should be solvable by eye

OpenStudy (konradzuse):

I could say u = cos(x) du = -sin(x)dx -du = sin(x)dx \[-\int\limits u^9 du\]

OpenStudy (turingtest):

exactly :)

OpenStudy (konradzuse):

\[\frac{-u^{10}}{10} +c\]

OpenStudy (konradzuse):

\[\frac{-cos^{10}}{10} +c\]

OpenStudy (konradzuse):

-cos(x)

OpenStudy (turingtest):

to make the cos part look better put a forward slash first \cos{whatever}\[\cos{whatever}\]

OpenStudy (konradzuse):

\[\cos(x)\]

OpenStudy (konradzuse):

NEXT QUESTION!!!